OpenStudy (anonymous):

please help!! Which numbers should be multiplied to obtain 183^2 - 139^2? 44 and 44 1936 and 1936 44 and 322 33350 and 33628

zepdrix (zepdrix):

Very strangely worded question.... I guess they want you to do the arithmetic, you'll get some big number... then they want to know which pair of numbers are factors of it.

zepdrix (zepdrix):

Were you able to do the first part? What is 183^2 - 139^2 ?

zepdrix (zepdrix):

Toss it into a calculator or something.

OpenStudy (anonymous):

okay one moment

OpenStudy (anonymous):

i get 14168

zepdrix (zepdrix):

Ok good! And which pair of numbers multiply to give you 14168?

OpenStudy (anonymous):

its c?

zepdrix (zepdrix):

Yay good job \c:/

OpenStudy (anonymous):

can you help me on another one?

OpenStudy (ranga):

\[\Large a^2 - b^2 = (a+b)(a-b) \\ \Large 183^2 - 139^2 = (183+139)(183-139) = 322*44 \]